OpenStudy (anonymous):

what are the coordinates of the terminal point determined by θ=7π/6?

OpenStudy (anonymous):

find \(\frac{7\pi}{6}\) on the unit circle on the last page of the attached cheat sheet and you will see them then i will show you how to find it if you can't use a cheat sheet

OpenStudy (anonymous):

oh thanks :) i found it, i'm just not sure how to use that

OpenStudy (anonymous):



OpenStudy (anonymous):

read off the coordinates

OpenStudy (anonymous):

they are the coordinates of the ordered pair you see at \(\frac{7\pi}{6}\) let me know what you see there

OpenStudy (anonymous):

i see -√(3)/2, 1/2

OpenStudy (anonymous):

then you win

OpenStudy (anonymous):

oh wait, they should both be negative right?

OpenStudy (anonymous):

yes, sorry i forgot to type the negative sign there

OpenStudy (anonymous):

ok good

OpenStudy (anonymous):

so is that the answer or do i have to do more? i'm confused.. sorry

OpenStudy (phi):

what part do you not yet understand ? 1) do you know you start reading angles from the x-axis, on the right side ?  2) 2 pi is all the way around. (If you convert 2 pi to degrees you will get 360 degrees) 3) half way is pi (or 180º have you ever heard someone say " he did a 180" meaning turned and went in the opposite direction ?